# Studying geometry of the ultraluminous X-ray pulsar Swift J0243.6+6124 using X-ray and optical polarimetry

### v1

@@ Acknowledgements.
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has been supported by the Academy of Finland grants 333112, 349144 and 355672 (JP, SST, AV), the German Academic Exchange Service (DAAD) travel grant 57525212 (VD, VFS), UKRI Stephen Hawking fellowship (AAM), Finnish Cultural Foundation grant (V. Kra.), Natural Sciences and Engineering Research Council of Canada (JH), and Deutsche Forschungsgemeinschaft (DFG) grant WE 1312/59-1 (VFS). I.L. was supported by the NASA Postdoctoral Program at the Marshall Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA. Nordita is supported in part by NordForsk.References
@@
- Heyl et al. (2024) Heyl, J., Doroshenko, V., González-Caniulef, D., et al. 2024, Nature Astronomy, in press, arXiv:2311.03667
@@
- Majumder et al. (2024) Majumder, S., Chatterjee, R., Jayasury, K. M., Das, S., & Nandi, A. 2024, arXiv e-prints, arXiv:2402.11602
@@
- Marin et al. (2024) Marin, F., Marinucci, A., Laurenti, M., et al. 2024, A&A, in press, arXiv:2403.02061
@@
- Veledina et al. (2024) Veledina, A., Muleri, F., Poutanen, J., et al. 2024, Nature Astronomy, in press, arXiv:2303.01174

### v2

@@ Acknowledgements.
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has been supported by the Academy of Finland grants 333112, 349144 and 355672 (JP, SST, AV), the German Academic Exchange Service (DAAD) travel grant 57525212 (VD, VFS), UKRI Stephen Hawking fellowship (AAM), Finnish Cultural Foundation grant (VKra), Natural Sciences and Engineering Research Council of Canada (JH), and Deutsche Forschungsgemeinschaft (DFG) grant WE 1312/59-1 (VFS). IL was supported by the NASA Postdoctoral Program at the Marshall Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA. Nordita is supported in part by NordForsk.References
@@
- Forsblom et al. (2024) Forsblom, S. V., Tsygankov, S. S., Poutanen, J., et al. 2024, A&A, in press, arXiv:2406.08988
@@
- Heyl et al. (2024) Heyl, J., Doroshenko, V., González-Caniulef, D., et al. 2024, Nature Astronomy, 8, 1047
@@
- Majumder et al. (2024) Majumder, S., Chatterjee, R., Jayasurya, K. M., Das, S., & Nandi, A. 2024, ApJ, 971, L21
@@
- Marin et al. (2024) Marin, F., Marinucci, A., Laurenti, M., et al. 2024, A&A, 689, A238
@@
- Veledina et al. (2024) Veledina, A., Muleri, F., Poutanen, J., et al. 2024, Nature Astronomy, 8, 1031